# Beyond the Surface: A Global-Scale Analysis of Visual Stereotypes in Text-to-Image Generation

### 12 Jan 2024 v1

@@ Abstract
Recent studies have highlighted the issue of stereotypical depictions for people of different identity groups in Text-to-Image (T2I) model generations. However, these existing approaches have several key limitations, including a noticeable lack of coverage of global identity groups in their evaluation, and the range of their associated stereotypes. Additionally, they often lack a critical distinction between inherently visual stereotypes, such as ‘underweight’ or ‘sombrero’, and culturally dependent stereotypes like ‘attractive’ or ‘terrorist’. In this work, we address these limitations with a multifaceted approach that leverages existing textual resources to ground our evaluation of geo-cultural stereotypes in the generated images from T2I models. We employ existing stereotype benchmarks to identify and evaluate visual stereotypes at global scale, spanning 135 nationality based identity groups. We demonstrate that stereotypical attributes are thrice as likely to be present in images of these identities as compared to other attributes. We further investigate how disparately offensive the depictions of generated images are for different nationalities. Finally, through a detailed case study, we reveal how the ‘default’ representations of all identity groups have a stereotypical appearance. Moreover, for the Global South, images across different attributes are visually similar, even when explicitly prompted otherwise. CONTENT WARNING: Some examples may contain offensive stereotypes.
@@ 1 Introduction
Figure 1: We identify ’visual’ stereotypes in the generated images of the identity group by grounding the evaluations in existing textual stereotype benchmarks. Yellow boxes denote annotated visual markers of known stereotypes associated with the identity group in the image. We use Stable Diffusion Rombach et al. (2022) to generate images and evaluate them using the stereotypes present in the SeeGULL dataset Jha et al. (2023).
Text-to-Image (T2I) models are increasingly used for generating visual content from textual descriptions, often for downstream tasks such as creative designing, advertising, marketing, and education. More often than not, these T2I models have been trained on large amounts of unverified data from the internet and tend to associate certain words or phrases (either positive or negative) with specific visual concepts. Although, their training enables these models to generate images from simple textual description, it also amplifies and reflects the stereotypes present in the society. Recent studies have highlighted how the generated images for certain identity groups are more stereotypical than others Bianchi et al. (2023a); Qadri et al. (2023); Ghosh and Caliskan (2023). While they notably demonstrate the biases in current state-of-the-art T2I models and how they reinforce existing stereotypes, there are several limitations to their approach.
There is a noticeable lack of coverage of identity groups in the evaluation of stereotypes of the generated images – in terms of the number of global identity groups, their associated stereotypes, and the volume of the images evaluated. This limitation leads to further under-representation of already marginalized minority groups. Moreover, since stereotypes are socially and culturally situated, some stereotypical attributes can be represented more objectively and reliably in an image than others. For example, representation of an ‘attractive’ person, or a ‘terrorist’, is highly subjective and culturally influenced. Existing works on stereotype evaluation often conflate these subjective attributes with more objective visual attributes (like ‘sombrero’, ‘underweight’, etc.,) which can be more objectively identified in an image. Moreover, the current evaluations of the images generated from T2I models blur problematic depictions with stereotypes, and are not necessarily grounded in existing stereotype resources, thereby, impacting the overall stereotype evaluation.
In this work, we address the above limitations by presenting a systematic large-scale cross-cultural evaluation of regional stereotypes present in the generated images from T2I models. We focus on objectively visual stereotypes and evaluate their prevalence in the visual depiction of 135 identity groups across the globe. More specifically, we ground our evaluations in an existing broad-coverage geo-cultural textual dataset, SeeGULL Jha et al. (2023) that incorporates stereotypes for 175 identity groups for the regional dimension. We also study how the default representation of an identity group relate to their stereotypical as well as non-stereotypcal representations. We use Stable Diffusion Rombach et al. (2022) as a representative T2I model for generating images and evaluate over 2000 images across different identity groups. Figure 1 provides a thematic representation of our approach. Although, our methodology uses Stable Diffusion and the SeeGULL repository as a reference, our overall approach is easily extensible to other textual datasets and T2I models. Our main contributions are as follows.
@@
  • We provide a nuanced approach for evaluating Text-to-Image generations for cross-cultural regional stereotypes at scale and with global coverage by leveraging existing resources from textual modality.
  • We make a critical distinction between objectively ‘visual’ and ‘culturally subjective’ stereotypes and identify a list of 385 ‘visual’ attributes that we will release along with their annotations.
@@
  • We introduce a broad-coverage image dataset [^1] For access to the dataset, please fill out the form at : https://forms.gle/3kqsry11D4gLKHQc7 with annotations of visual markers of stereotypes present in the images representing different identity groups.
@@
  • We analyze the representations of several identity groups and demonstrate how T2I models disproportionately lean towards stereotypical representations of identity groups, even when explicitly prompted otherwise.
@@ 2 Related Work
Stable Bias Luccioni et al. (2023) identifies social biases in Text-to-Image systems by analyzing variations in generated images based on gender, ethnicity markers, and professions. They evaluate three popular T2I systems (DALL-E, Stable Diffusion V1.4 and 2), and demonstrate the under-representation of marginalized identities in the generated images. Bianchi et al. (2023b) reveal how T2I models perpetuate dangerous stereotypes related to traits, occupations, and objects using different prompts. They also demonstrate how user efforts to counteract the biases through different prompting strategies prove unsuccessful. Qadri et al. (2023) explore the cultural limitations of T2I models, specifically in South Asian context, and use community-centered perspective to demonstrate how T2I models use an ‘outsiders gaze’ when generating images for South Asian identity groups. Hao et al. (2023) provide a theoretical framework for content moderation for T2I models in their work and propose ways to empirically measure and quantify the harms associated with the generative models. Garcia et al. (2023) partially annotate Google Conceptual Captions dataset with four demographic and two contextual attributes, and demonstrate the demographic biases in image captioning, text-to-image CLIP embeddings, and images generated from T2I models. Naik and Nushi (2023) quantify social biases in occupations, personality traits, and everyday situations, and show that the images generated for default location-neutral prompts resemble those from the United States and Germany. They also demonstrate that not all biases can be mitigated by adding specifications in the input prompt. Basu et al. (2023) quantify the geographical representativeness of the images generated from T2I models. Wang et al. (2023) design a text-to-image association test to measure the implicit stereotypes between concepts and valence, and and their presence in the generated images. Recently, Gustafson et al. (2023) released the FACET dataset containing annotations about person-related attributes and objects such as skin-tine, and measure performance disparity in image classification, object detection and segmentation tasks. Although, there has been a lot of interest in stereotype evaluation recently, we are the first ones to provide a large-scale evaluation study of ‘regional’ stereotypes in T2I models by leveraging existing resources from textual modality. Moreover, we explicitly focus on identifying stereotypes that are primarily visual in nature to reduce the subjectivity in stereotype evaluation.
@@
There have been several recent efforts to build broad-coverage cross-cultural stereotype resources at scale within textual modalities. For example, StereoSet Nadeem et al. (2021) and CrowS-Pairs Nangia et al. (2020) are stereotype datasets for detecting stereotypes in NLP prediction and classification-based tasks. SeeGULL Jha et al. (2023) is another such stereotype repository which has a broader coverage of global stereotypes containing close to 7000 stereotypes about identity groups spanning 178 countries across 8 different geo-political regions across 6 continents. It also includes stereotypes for state-level identities within the US and India. In this work, we leverage resources from textual modalities, such as SeeGULL, to conduct a comprehensive analysis of stereotypes prevalent in the images generated from Text-to-Image (T2I) models.
3 A Nuanced Analysis of Stereotypes in Images
Stereotypes and their visual depictions are culturally influenced, highly subjective, and vary across regions. For example, in North America, perceptions such as Australians being ‘attractive’, Northern Irish as ‘terrorists’, and Chinese as ‘intelligent’ are quite common Jha et al. (2023). However, these stereotypes are not as common in other countries Jha et al. (2023). The stereotypical attributes lack inherent visual characteristics, and their portrayal is heavily influenced by cultural perceptions. The depiction of a ‘terrorist’, an ‘attractive’ person, or an ‘intelligent’ person, therefore, can vary significantly based on the geographical and cultural context. In contrast, attributes like ‘sombrero’ (often associated with Mexicans), ‘skinny’, ‘underweight’, ‘black’ (stereotypically associated with Sudanese), etc., possess distinct visual markers, allowing a more objective depiction in images (Figure 1). We note here that the terms ‘black’, ‘white’ etc. in this work, as identified in images do not imply race, and instead are about appearance, or appearance-based observed race Hanna et al. (2020), or even the colors themselves.
To ensure a more reliable evaluation of the generated images from Text-to-Image Models for stereotypes, it is crucial to mitigate this subjectivity in visual representation by making a clear distinction between visually identifiable and culturally subjective stereotypes. Therefore, we follow a nuanced two-step approach for evaluating stereotypes in T2I models. As a first step, we answer the fundamental question of which stereotypes can be objectively represented in an image. We identify inherently ‘visual’ attributes which we subsequently use in the second step to detect stereotypes in the generated images by (i) conducting a large-scale annotation study, and (ii) using automated methods. These steps are described in more detail below.
@@ 3.1 Identifying Visual Stereotypes
As propounded above, we first undertake a systematic annotation task for annotating ‘visual’ attributes to make a clear distinction between visual and non-visual/subjective attributes. We use an existing stereotype resource in textual modality, SeeGULL Jha et al. (2023), as a reference to ground our evaluations. SeeGULL comprises of (identity, attribute) pairs, where the ‘identity’ term denotes the global identity groups, and ‘attribute’ denotes the associated descriptive attribute terms (adjective/adjective phrase, or a noun/noun phrase), such as (Mexicans, sombrero), (Indians, dark), and (English, stiff upper lip). The dataset has 1994 unique attributes – both visual and non-visual. We conduct an annotation task to separate out the visual attributes from the non-visual ones.
Annotating Visual Nature of Attributes
@@
The annotators are presented with an attribute and a statement of the form of ‘The attribute [attr] can be visually depicted in an image.’ The attribute presented in the sentence is a list of all unique attributes (noun/noun phrases, and adjective/adjective phrases) from SeeGULL. The annotators are required to assign a Likert scale rating to each attribute indicating the extent to which they agree or disagree with the above statement w.r.t. to the given attribute term. The Likert Scale labels were as follows:
  • Strongly Agree: When the attribute can be explicitly identified within an image, like objects, colors, or similar visual elements (e.g., ‘hat,’ ‘sombrero,’ ‘short’).
  • Agree: When the attribute can be deduced from visual cues, albeit not explicitly depicted in the image (e.g., ‘fashionable,’ ‘poor,’ ‘impoverished’).
  • Disagree: When the attribute is challenging to detect visually but may be inferred from visual cues in specific contexts.
  • Strongly Disagree: When the attributes cannot be inferred visually, either explicitly or through visual cues, such as ‘kind,’ ‘talkative,’ ‘warmhearted,’ and the like.
  • Unsure: When annotators are uncertain about the attribute’s visual nature.
Annotator Recruitment
@@
Since the task was in English, we recruited annotators proficient in English reading and writing. The raters were based out of United States. We get annotations for each attribute from 3 annotators that identify with different geographical orgin identities - Asian, European, and North American. Annotators were professional data labelers and were compensated at rates above the prevalent market rates, and respecting the local regulations regarding minimum wage in their respective countries.
Identifying Associated Visual Stereotypes
@@
Using the identified ‘visual’ attributes (please refer to the Appendix 8.1 for results), we filter out the associated ‘visual stereotypes’ in SeeGULL. Figure  2 shows the global distribution of visual attributes and contains examples of visual stereotypes associated with different countries. We observe that Australians (63) have the highest number of visual attributes associated with them followed by Mexicans (46), Indians (34), New Zealand (31), Ethiopians (27), and Japanese (20). The numbers in bracket indicate the number of visual attributes associated with the identity group in SeeGULL. Around 30 identities, including Omani, Ukranian, Swiss, Canadian, Mongolian, etc., just have a single visual attribute associated with them. We use these visual stereotypes for identifying their prevalence in images in the next step.
@@
Figure 2: Global distribution of visual stereotypes across countries. The depth of the color indicates the number of visual stereotypes. A few examples of visual stereotypes of some of the countries have been shown in the figure.
@@ 3.2 Detecting Visual Stereotypes in Text-to-Image Generation
More often than not, stereotypes reflected in the images generated using T2I models are implicit associations learned by the models about an identity group during their training. Evaluating these images for the presence of stereotypes is essential so as to not amplify or reinforce the existing bias in the society which can lead to further discrimination of marginalized groups. As a second step to identifying visual stereotypes, we conduct a large-scale annotation task to detect prevalent ‘visual’ stereotypes in the generated images of 135 identity groups across the globe. The scale of our evaluation is significantly larger than the prior work – in terms of image volume, the global coverage of identity groups, and their associated stereotypes.
Generating Images using T2I Models
[…]
Annotating Stereotypes in Images
@@
The annotation task aims to examine whether well-known ‘visual’ stereotypes about various identity groups are reflected in the images generated by Text-to-Image models. We generate 15 images per identity group and present it to the annotators. Each image is accompanied by two sets of attributes: (i) a set of visual attributes stereotypically associated with the identity group (Figure 2), and (ii) an equal number of randomly selected visual non-stereotypical attributes, i.e., attributes that are not stereotypically associated with the identity group. We combine the two sets and randomly show 5 attributes in succession alongside each image, until every attribute has been covered. An additional option of ‘None of the above’ is also displayed with each set. The annotators are asked to select all attribute(s) that they believe are ‘visually’ depicted in the image. Additionally, they are also asked to draw bounding boxes to highlight specific regions, objects, or other indicators that support their selection of the visual attribute within the image. If annotators believe that no attributes are visually represented, they can choose ‘None of the above.’
Annotator Recruitment
Since the task was in English, we recruited annotators proficient in English reading and writing. The raters were based out of United States but identified with different nationalities. We get 3 annotations for each attribute-image pair. Annotators were professional data labelers and were compensated at rates above the prevalent market rates, and respecting the local regulations regarding minimum wage in their respective countries.
Automatically Identifying Stereotypes in Images
Annotating stereotypes in images at scale can often times be resource-intensive and time-consuming. Therefore, in addition to annotating stereotypes in images, we also investigate the feasibility of using automated techniques for their detection in images. We use the same set of 15 images per identity group and use an already existing image captioning model, CLIP (Contrastive Language-Image Pre-training) to get top 50 cations or text description of each image.
@@
Building on the already identified visual stereotypes (Section 3), we conduct an analysis of the captions generated for the images of each of the 135 identity groups. We identify the stereotypes in the generated captions by performing a string match. To further understand how uniquely a stereotypical attribute $attr_{s}$ is present in the description of the images of the identity group, we compute a salience score of the attributes w.r.t. the identity group $S(attr_{s},id)$ by using a modified tf-idf metric as follows: $S(attr_{s},id)=tf(attr_{s},id)\cdot idf(attr_{s},C)$ The function $tf(attr_{s},id)$ represents the smoothed relative frequency of attribute $attr_{s}$ for the identity group $id$; and the function $idf(attr_{s},C)$ denotes the inverse document frequency of the attribute term $attr_{s}$ in all the captions ‘C’, reflecting the importance of the attribute across all the captions. A higher salience score of an attribute, indicated a more unique association of the attribute with the identity group. For each identity group, we then extract the most salient visual stereotypes associated with it as present in the captions.
4 Analyzing Stereotypical Depictions in Generated Images

### 22 Feb 2024 v2

# ViSAGe: A Global-Scale Analysis of Visual Stereotypes in Text-to-Image Generation

@@ Abstract
Recent studies have shown that Text-to-Image (T2I) model generations can reflect social stereotypes present in the real world. However, existing approaches for evaluating stereotypes have a noticeable lack of coverage of global identity groups and their associated stereotypes. To address this gap, we introduce the ViSAGe (Visual Stereotypes Around the Globe) dataset to enable evaluation of known nationality-based stereotypes in T2I models, across 135 nationalities. We enrich an existing textual stereotype resource by distinguishing between stereotypical associations that are more likely to have visual depictions, such as ‘sombrero’, from those that are less visually concrete, such as ‘attractive’. We demonstrate ViSAGe’s utility through a multi-faceted evaluation of T2I generations. First, we show that stereotypical attributes in ViSAGe are thrice as likely to be present in generated images of corresponding identities as compared to other attributes, and that the offensiveness of these depictions is especially higher for identities from Africa, South America, and South East Asia. Second, we assess the stereotypical pull of visual depictions of identity groups, which reveals how the ‘default’ representations of all identity groups in ViSAGe have a pull towards stereotypical depictions, and that this pull is even more prominent for identity groups from the Global South. CONTENT WARNING: Some examples contain offensive stereotypes.
@@ 1 Introduction
Figure 1: We identify ‘visual’ stereotypes in the generated images of the identity group by grounding the evaluations in existing textual stereotype benchmarks. Yellow boxes denote annotated visual markers of known stereotypes associated with the identity group in the image. We use Stable Diffusion Rombach et al. (2022) to generate images and evaluate them using the stereotypes present in the SeeGULL dataset Jha et al. (2023).
Text-to-Image (T2I) models are increasingly being used to generate visual content from textual descriptions Rombach et al. (2022); Ramesh et al. (2021), enabling downstream tasks such as creative design, advertising, marketing, and education. More often than not, these T2I models have been trained on large amounts of web-scale data with minimal curation, if any. As a result, these models often reflect and propagate stereotypes about identities and cultures present in the data Bianchi et al. (2023); Ghosh and Caliskan (2023); Luccioni et al. (2023).
While these studies provide crucial evidence for fairness failures of T2I generations, they have a noticeable lack of coverage of global identity groups and their associated known stereotypes. Some studies offer qualitative insights across different identity groups but focus on only a limited set of identities and known stereotypes Bianchi et al. (2023); Qadri et al. (2023). On the other hand, existing larger scale studies focus primarily on stereotypes in the US or Western society Luccioni et al. (2023), and lack global identities and contexts. Given the widespread adoption of these models in global contexts, it is imperative to build extensive, stratified evaluations that cover broader identity groups to prevent further under-representation of already marginalized groups.
We address the above challenges by presenting a systematic, large-scale, cross-cultural evaluation of regional stereotypes present in the generated images from T2I models. Specifically, we ground our evaluations in an existing geo-culturally broad-coverage, textual resource, SeeGULL Jha et al. (2023) that incorporates societal stereotypes for 175 nationality based identity groups. This grounding in existing social stereotype resources aids the critical distinction between spurious correlations in models and stereotypical tendencies which are necessary for model safety interventions Blodgett et al. (2021); Selvam et al. (2023). We enrich this data with human annotations to incorporate the notion of which stereotypes are more likely to be visually depicted versus those that cannot be. Building on this societal grounding, we conduct a study to investigate the extent of stereotypical and offensive depictions of different nationalities in a scalable manner. We also study how the default representation of a nationality relates to their stereotypical as well as non-stereotypical representations. Figure 1 provides a thematic representation of our approach. Our main contributions are as follows.
@@
  • We evaluate Text-to-Image generations for cross-cultural regional stereotypes at scale and with global coverage by leveraging existing resources from textual modality.
@@
  • We publicly release the dataset ViSAGe : Visual Stereotypes Around the Globe where we make a critical distinction between ‘visual’ and ‘non-visual’ stereotypes in images. We identify a list of 385 visual attributes, and also introduce a broad-coverage image dataset with annotations of visual markers of stereotypes present in 40,057 image-attribute pairs, representing different identity groups.  [^1] For access to the dataset, please fill out the form at : https://forms.gle/3kqsry11D4gLKHQc7
@@
  • We analyze the default representations of identity groups and demonstrate how T2I models disproportionately lean towards their stereotypical representations, even when explicitly prompted otherwise.
@@ 2 Related Work
Cho et al. (2023) show that T2I models reflect specific gender/skin tone biases from training data. Zhang et al. (2023) study gender presentation in T2I models. Bianchi et al. (2023) highlight the presence of stereotypes by using prompts containing attributes and associating the generated visual features with demographic groups. Stable Bias Luccioni et al. (2023) prompts T2I models with a combination of ethnicity/gender and profession and evaluates profession-based stereotypes. Basu et al. (2023) use location-based prompts to quantify geographical representativeness of the generated images. Qadri et al. (2023) identify harmful stereotypes and an ‘outsiders gaze’ in image generation in the South Asian context. We prompt T2I models with ‘identity groups’ around the globe to study the prevalent regional stereotypes in their default visual representation. We identify ‘visual’ stereotypes, related to the concept of ‘imageable’ synsets Yang et al. (2020), to conduct this analysis at scale.
Several efforts have also been made to quantify the harms associated with generative models. Hao et al. (2023) propose a theoretical framework for content moderation and its empirical measurement. Naik and Nushi (2023) and Wang et al. (2023) quantify social biases in T2I generated images. Garcia et al. (2023) and Gustafson et al. (2023) highlight demographic biases in image captioning and classification tasks. We investigate the feasibility of automatically identifying visual stereotypes in image generations, and demonstrate their offensiveness.
@@
StereoSet Nadeem et al. (2021) and CrowS-Pairs Nangia et al. (2020) are datasets used for detecting stereotypes in NLP prediction and classification-based tasks. SeeGULL Jha et al. (2023) is a stereotype repository with a broad coverage of global stereotypes – containing close to 7000 stereotypes for identity groups spanning 178 countries across 8 different geo-political regions across 6 continents. It comprises of (identity, attribute) pairs, where ‘identity’ denotes global identity groups, and ‘attribute’ denotes the associated descriptive stereotypical adjective/adjective phrase, or a noun/noun phrase, such as (Mexicans, sombrero), (Germans, practical), and (Japanese, polite). In this work, we leverage resources from textual modalities, such as SeeGULL, to ground our evaluation of prevalent stereotypes in the generated images.
3 Our Approach
Some stereotypes can be recognized in terms of clearly visual attributes (e.g.: ‘Mexicans, sombreros’) whereas other stereotypes are defined through non-visual characteristics such as social constructs, and adjectives (e.g.: ‘Chinese, intelligent’). While the latter may have some visual markers, they are difficult to represent accurately in images. Therefore, we follow a nuanced two-step approach. As a first step, we answer the fundamental question of which stereotypes can be objectively represented in an image. We identify inherently ‘visual’ attributes which we subsequently use in the second step to detect stereotypes in the generated images by (i) conducting a large-scale annotation study, and (ii) using automated methods.
@@ 3.1 Identifying Visual Stereotypes
To mitigate subjectivity in visual representation and ensure a more reliable evaluation of the generated images from T2I Models for stereotypes, it is crucial to make a distinction between clearly identifiable ‘visual’ stereotypes and difficult to represent ‘non-visual’ stereotypes. We undertake a systematic annotation task to differentiate the two. We use an existing stereotype resource in textual modality, SeeGULL Jha et al. (2023), as a reference to ground our evaluations. The dataset has 1994 unique attributes – both ‘visual’ and ‘non-visual. We deduce visual attributes and then map it back to the identity terms in SeeGULL to obtain visual stereotypes.
Annotating Visual Attributes
@@
The annotators are presented with an attribute and a statement of the form of ‘The attribute [attr] can be visually depicted in an image.’ The attribute presented in the sentence is selected from a list of all unique attributes (noun/noun phrases, and adjective/adjective phrases) from SeeGULL. The annotators are required to assign a Likert scale rating to each attribute indicating the extent to which they agree or disagree with the above statement w.r.t. to the given attribute term. We note that the visual nature of attributes lie on a spectrum. While there may be clearly defined attributes on either extremes, the ‘visual’ and ‘non-visual’ attributes in the middle are more subjective. For our subsequent analysis, we exclude all attributes where any annotator expressed uncertainty or disagreement regarding the visual nature. Consequently, we deem terms where all annotators at least agreed about their visual nature, as ‘visual’ resulting in a selection of 385 out of the original 1994 attributes. Please refer to the Appendix 7.1.1 for details.
@@
We recruited annotators based out of the United States, and proficient in English reading and writing. For each attribute, we get annotations from 3 annotators that identify with different geographical origin identities - Asian, European, and North American. Annotators were professional data labelers and were compensated at rates above the prevalent market rates.
Mapping to Visual Stereotypes
@@
Using the identified visual attributes (Appendix 7.1.1), we filter out the associated ‘visual stereotypes’ in SeeGULL. Figure  2 shows the global distribution of visual attributes. It also contains examples of visual stereotypes associated with different countries. We observe that around 30 identities, including Omani, Ukranian, Swiss, Canadian, Mongolian, etc., just have a single visual attribute associated with them. Australians (63) have the highest number of visual attributes associated with them followed by Mexicans (46), Indians (34), New Zealand (31), Ethiopians (27), and Japanese (20). The numbers in bracket indicate the number of visual attributes associated with the identity group in SeeGULL. We use these visual stereotypes for identifying their prevalence in images in the next step.
@@
Figure 2: Global distribution of visual stereotypes across countries. Depth of the color indicates the number of visual stereotypes. A few examples of visual stereotypes of some countries are shown in the figure.
@@ 3.2 Detecting Visual Stereotypes in Text-to-Image Generation
Evaluating generated images for the presence of stereotypes is essential to not amplify or reinforce existing biases in the society. We conduct a large-scale annotation task to detect prevalent visual stereotypes in the generated images of 135 identity groups across the globe. The scale of our evaluation is significantly larger than the prior work – in terms of the global coverage of identity groups, and their associated stereotypes.
We use Stable Diffusion-v1.4 Rombach et al. (2022) as the base Text-to-Image model, denoted as $M$, and generate images representative of various identity groups by defining three distinct prompts as follows: (i) $P_{1}({id})=\textnormal{`a photo of $id$ person'}$’, (ii) $P_{2}({id})=\textnormal{`a portrait of $id$ person'}$’, and (iii) $P_{3}({id})=\textnormal{`an $id$ person'}$’, where ${id}\in\mathbb{X}$ such that $\mathbb{X}$ is the set encompassing identity terms extracted from the SeeGULL dataset. For each prompt $P_{i}(i=1,2,3)$ and each identity term ${id}$, we generate 5 images, resulting in an output set of images $I=\{I_{i,1},I_{i,2},\ldots,I_{i,5}\}$ for a combination of prompt and identity term. The image generation process can be denoted as: $I_{i,j}=M(P_{i}({id})),$ where $j=(1,2,\ldots,5)$ is the generated images for a given prompt $P_{i}$ and identity term $id$.
Detecting Stereotypes through Human Annotations
@@
The annotation task aims to examine whether well-known visual stereotypes about identity groups are reflected in the generated images. We generate 15 images per identity group and present it to the annotators. Each image is accompanied by two sets of attributes: (i) a set of visual attributes stereotypically associated with the identity group (Figure 2), and (ii) an equal number of randomly selected visual non-stereotypical attributes, i.e., attributes that are not stereotypically associated with the identity group. We combine the two sets and randomly show 5 attributes in succession alongside each image, until every attribute has been covered. An additional option of ‘None of the above’ is also displayed with each set. The annotators are asked to select all attribute(s) that they believe are visually depicted in the image. Additionally, they are also asked to draw bounding boxes to highlight specific regions, objects, or other indicators that support their selection of the visual attribute within the image. If annotators believe that no attributes are visually represented, they can choose ‘None of the above.’ Overall, we get annotations for 2,025 identity-image pairs and 40,057 image-attribute pairs. Similar to  3.1, we recruited annotators proficient in English reading and writing, residing in United States but identifying with different nationalities. We get annotations for each attribute-image pair from 3 annotators.
Detecting Stereotypes through Automated Methods
@@
Annotating stereotypes in images at scale can often times be resource-intensive and time-consuming. Therefore, we also investigate the feasibility of using automated techniques for stereotype detection in images. Using an already existing image captioning model, CLIP (Contrastive Language-Image Pre-training), we get top 50 captions for each image for the same set of the 15 images per identity group previously generated. Building on the already identified visual stereotypes, we identify the stereotypes in captions by performing a string match. To further understand how uniquely a stereotypical attribute $attr_{s}$ is present in the caption of the images of an identity group, we compute a salience score of the attributes w.r.t. the identity group $S(attr_{s},id)$. We use a modified tf-idf metric as follows: $S(attr_{s},id)=tf(attr_{s},id)\cdot idf(attr_{s},C)$. The function $tf(attr_{s},id)$ represents the smoothed relative frequency of attribute $attr_{s}$ for the identity group $id$; and the function $idf(attr_{s},C)$ denotes the inverse document frequency of the attribute term $attr_{s}$ in all the captions ‘C’, reflecting the importance of the attribute across all the captions. A higher salience score of an attribute, indicates a more unique association of the attribute with the identity group. For each identity group, we then extract the most salient associated visual stereotypes as present in the captions.
4 Study 1: Stereotypical Depictions